An economy is described by the solow model, it has he following production function: Y= F(K,EL) K5 (EL) 0.5 E grows at rate g; L grows at rate n ; depreciation rate is ô. Savings rate is a constant s 1- We will fill in the model (in terms of Y, s) C= (in terms of Y, s Y = (in terms of C and I only) 2- This is the first year of our country's founding, the country is called NewNew England, NNE for short. it has a 100 people (L) and Capital (K). Every person in NNE saves 25% of their income. And existing capital depreciates at a rate of 7%. E= 1, both g and n equal zero. a 100 units of L T Y1 C. I 8 K Kta 0 3- We want to describe the economy per effective worker so, y=Y/EL ,k=K/EL ,c=C/EL ,i= 1/EL Derive the per capita production function y (hint, y will be expressed in terms of k) show your work SC
y ? (in terms of y, s) (in terms of y, s) i = 4- Capital per person changes according to this law k+i-(6+n+ g) k kt+1 can you very briefly explain the intuition behind this equation? In the steady state, k= k+1 , derive the steady state k' equation in terms of s, 8, n,g the equations for i, y and the law of motion) (hint, use k' ?
Answer #1

yeFIKEL]-KaLEL. 1 를 - 눈= 8 A savings ale & dpuasin Cd) ate n It knoun that y=C+I& aloo (Z) Y-C+S y=C+ S y: ct69 opotion of y

(11 Ct C1 A) 1 = (1-.25) x 108.62= 81.465 C1= 81.47 1 2 I1: 1 27.15 13 ,25X108.62 SKI=.07 X 11g 8.26 SK1K2KI + T1-8 KI 11 K2-

(-s)Y-6) --4)k EL EL U-s)KOS 0.5 34 REtl =ke +it- C8tnt kE intuitiom- (Retl-kt) - Ak it-(8+nt9) RE eguation depiets the cnmge

as Rt kG+1 Jue dtrady state is olafined te -(Stntg)ke i ke-keti (REHI-RE) 0 i6-(8tntg) ke it = (8tntz)kt L It =s YE EL EL Rt